OpenStudy (anonymous):

The ellipse (x^2)/25 + (y^2)/36 = 1 , can be drawn with parametric equations. If x = r cos(t) what's r and y??

OpenStudy (anonymous):

Fruitless answerer: x=5*cos(t) y=+-(6/5)*sqrt(25-x^2) thank you

OpenStudy (m):

do you know where x = rcost comes from? think about x^2 + y^2 = r^2 and cos^2t + sin^2t = 1

OpenStudy (anonymous):

r is 5 x = 5 * cos(t) so y = +-(6/5)*sqrt(25-(5*cos(t))^2)

OpenStudy (anonymous):

x=5cos(t) y=5sin(t)

OpenStudy (anonymous):

sont know why frutless has given a mess for the value of y, it should be expressed in terms of the parameter as well